# Trapped-ion quantum simulation of the Fermi-Hubbard model as a lattice gauge theory using hardware-aware native gates

### v1

@@ III.3 Determining simplifiable subroutines
The objective is to create a parametrizable circuit (in number of gates and overall depth) $\hat{C}_{A}$, known as an ansatz, that satisfies $\hat{C}_{A}\approx\hat{C}$. Our ansatz is shown in Fig. 2, and is comprised of single qubit $R_{X}(\theta)$ and $R_{Y}{\theta}$ gates, and a two qubit $XX(\theta)$ gate.
@@
In each iteration of gradient-based optimization (GBO), gate angles are generated by an optimizer. Then, using the matrix representation of $\hat{C}$ and $\hat{C}_{A}$, the similarity between the two circuits is given by $1-\mathcal{F}$, where
@@
The efficacy of GBO in converging to a solution is limited by local minima, barren plateaus, initial conditions, noise when computing the gradient, and the number of tunable parameters. A larger depth $l$ results in more parameters to optimize over and the introduction of further minima and plateaus in the optimization problem. Therefore, an optimizer that is consistent in performance and resilient to these challenges is critical to finding gate angles that satisfy $\hat{C}\approx\hat{C}_{A}$.
We present iteratively pre-conditioned gradient descent (IPG) [34, 51] and find suitable gate angles to surmount these challenges. IPG is a noise-resilient, higher-order algorithm that has shown promising gains in convergence speed for classical optimizations, converging locally at a linear rate for convex problems and superlinearly when the solution is unique. We have previously shown the utility of the IPG algorithm on quantum circuit optimization [52] whereby faster convergence was observed over other optimizers like gradient descent and Adam [53]. Nevertheless, the exponential growth of the Hilbert space dimension with the number of qubits makes brute-force IPG infeasible for classically optimizing the entire quantum circuit. Instead, here we identified a 3-qubit unitary operator $\hat{C}$ which repeatedly appears at each Trotter step and optimize this subcircuit such that the size of the circuit to be optimized stays fixed independent of the number of sites $N$ or qubit count $3N$. This leverages the model’s translational symmetry discussed in Sec. III.1 and the all-to-all connectivity of the trapped-ion QPU which makes implementing periodic boundary conditions connecting $i=1$ to $i=N$ straightforward.
IV Iteratively Pre-Conditioned Gradient Descent
This section introduces the Iteratively Preconditioned Gradient (IPG) descent algorithm, which computes a minimizer of the cost function $f:\mathbb{R}^{d}\to\mathbb{R}$. Formally, the goal is to compute a gate-angle vector $x_{*}\in\mathbb{R}^{d}$ such that
@@ IV Iteratively Pre-Conditioned Gradient Descent
When the cost function is non-convex, instead of searching for a globally optimal solution, a more meaningful and attainable goal is to find a local minima of the cost function $f$, defined as a point such that, for some $\epsilon>0$ every point satisfying $|x^{*}-x|_{2}<\epsilon$ must also satisfy $f(x)>f(x^{*})$, where $\nabla f(x)\in\mathbb{R}^{d}$ denotes the gradient of $f$ at $x\in\mathbb{R}^{d}$.
Built upon the prototypical gradient-descent (GD) algorithm [54], several accelerated gradient algorithms have been proposed for solving (14[55, 56, 57, 58]. Amongst them, some notable algorithms are Nesterov’s accelerated gradient-descent (NAG) [55], the heavy-ball method (HBM) [56], and adaptive momentum (Adam) [53]. Although these methods improve GD’s convergence rate, they converge at a sublinear rate [59, 60] for general nonconvex costs.
By contrast, Newton’s method [61] offers another method of accelerating convergence by using the Hessian of $f$. Specifically, Newton’s method uses preconditioning, which pre-multiplies the gradient with the inverse Hessian matrix at every iteration, resulting in local quadratic convergence rates [61] for strongly convex objectives. Despite its faster convergence rate, there are several issues in Newton’s method. (i) For empirical risk minimization, the per-iteration computational cost of Newton’s Method is $\text{O}(nd^{2}+d^{3})$. (ii) Secondly, the accelerated convergence of Newton’s method is guaranteed only if $f$ is strongly convex, an unlikely condition in quantum circuit optimization. (iii) It involves computing a matrix inverse at every iteration, which is highly unstable against process noise, such as qubit decoherence and gate infidelity. While Newton’s method offers superior convergence rates, these issues ultimately mean it is ill-suited for use in this setting.
@@
Algorithm 1 Iteratively Preconditioned Gradient Descent Algorithm
1:Starting at $x_{0}\in\mathbf{X}$, objective function $f$, initial preconditioner estimate $K_{0}$, stabilization parameter $\beta\geq 0$, initial parameter descent rate $\delta>0$
2:for $t=0$ to $n$ do
3:     $x_{t+1}\leftarrow x_{t}-\delta_{t}K_{t}\nabla f_{t}$
4:     $\delta\leftarrow 1/\left(\lambda_{\rm max}(\nabla^{2}f_{t})+\beta\right)$
5:     $K_{t+1}\leftarrow K_{t}-\delta(\nabla^{2}f_{t}K_{t}+\beta K_{t}-I)$
6:end for
The IPG algorithm is structured similarly to Newton’s method in that both approaches rely on preconditioning the gradient to accelerate convergence. Still, they differ in the way the preconditioner is constructed. Specifically, rather than explicitly taking the inverse Hessian like Newton’s method, IPG finds an approximate inverse with offset eigenvalues to help maintain stability. The notable difference with quasi-Newton methods, such as BFGS [61], which also construct an approximation of inverse Hessian, is that the preconditioner in IPG is not required to be symmetric and positive definite, a condition likely to be violated when implemented with noisy quantum gates. Next, we present the steps of Algorithm 1.
Initialization: Before starting the iterations, an initial estimate $x_{0}$ and a pre-conditioner matrix $K_{0}$ is chosen from $\mathbb{R}^{d}$ and $\mathbb{R}^{d\times d}$, respectively. Further, three sequences of non-negative scalar parameters $\{\alpha_{t},\beta_{t},\delta_{t},t\geq 0\}$ are chosen, such that $\delta_{t}\leq 1$, $\beta_{t}>-\lambda_{\min}(H_{t})$, and $\alpha_{t}<1/(\lambda_{\max}(H_{t})+\beta_{t})$. Here, $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ respectively denote the smallest and the largest eigenvalue of a square matrix.
Steps in each iteration $\bm{t}$: For each iteration $t\geq 0$, we let $f_{t}=f(x_{t})$, $g_{t}=\nabla f(x_{t})$, and $H_{t}=\nabla^{2}f(x_{t})$ respectively denote the value of the cost function $f$, its gradient vector, and the Hessian matrix evaluated at the current estimate $x_{t}$. Let $I$ denote the $(d\times d)$-dimensional identity matrix. In each iteration $t$, the algorithm comprises two steps. In Step 1, the estimate $x_{t}$ is updated to $x_{t+1}$ such that
$x_{t+1}$ $=x_{t}-\delta_{t}K_{t}\nabla f_{t}.$ (15)
In Step 2, the pre-conditioner matrix $K_{t}$ is updated to $K_{t+1}$ using second-order information $H_{t}$:
$K_{t+1}$ $=K_{t}-\alpha_{t}\left(\left(\nabla^{2}f_{t}+\beta_{t}I\right)K_{t}-I\right).$ (16)
In deterministic settings, the convergence analysis of the IPG algorithm can be found in [62]. In the presence of noise, $x_{t}$ in the IPG algorithm converges to a neighborhood of a stationary point $x_{st}$ [63]. Empirically, for solving classical optimization problems, IPG requires fewer iterations to reach the desired accuracy and obtains a smaller steady-state error than prominent gradient-based optimizers [64].
The aforementioned IPG algorithm has recently been utilized for optimizing gate-based quantum circuits [52]. Specifically, IPG showed faster convergence to a higher fidelity for quantum state preparation and quantum algorithmic subroutines, such as approximating the full unitary matrix of a quantum Fourier transform and quantum state preparation. However, these results with IPG in [52] were generalized for most gate-based QPUs. Particularly, the hardware-agnostic approach in [52] was insufficient to benchmark the optimized circuits on the Aria-1 QPU. In this paper, we investigate the performance of IPG on optimizing hardware-aware (IonQ’s trapped-ion) ansatz presented in Section III.
V Optimizing the ansatz
We now optimize the gate angles in $\hat{C}_{A}$ at a depth of $l=3$ layers (Fig. 2a) using IPG, and benchmark IPG performance to L-BFGS and Adam, two optimizers commonly used for quantum circuit optimization. Cost is defined as $1-\mathcal{F}$, where lower cost indicates better agreement between $\hat{C}_{A}$ and $\hat{C}$.
Fig. 2 shows results for circuit optimization for each optimizer. IPG converges to the lowest cost in the fewest iterations, approximately four orders of magnitude lower than L-BFGS.
To verify that the converged gate angles of each optimizer satisfy $\hat{C}_{A}\approx\hat{C}$, we apply 10000 random input states |$\ket{\psi_{\rm in}}$ on both circuits and take the average of the state infidelity as
@@ V Optimizing the ansatz
for each resulting output state |$\ket{\psi_{C_{A,[i]}}}=\hat{C}_{A,[i]}\ket{\psi_{\rm in}}$. In doing so we find that only the gate angles generated by IPG are suitable enough to represent $\hat{C}$.
@@
After constructing the overall Trotter step using $\hat{C}_{A}$ with angles converged to by IPG, we verify that the reconstructed trotter step is the same as the original, using the convergence method outlined above. The average difference between states generated by a single parameterized vs explicit trotter step is of order $10^{-4}$, sufficiently low to consider both circuits approximately equal.
@@
Figure 2: a) Modified ansatz $\hat{C}_{A}$ from [30] used for gradient based optimization to parametrize $\hat{C}$. Each layer is comprised of $R_{X}(\theta_{1})$, $R_{y}(\theta_{2})$, $XX(\theta_{3})$ on qubits $S_{i,\sigma}$, $B_{i,i+1}$, and $S_{i+1,\sigma}$ respectively. Three layers are used with a final $R_{X}(\theta)$ and $R_{Y}(\theta)$ rotation on each qubit. b) Result of GBO for 30 parameters using IPG, Adam, and L-BFGS. Best cost history over three trials and 128 iterations for each optimizer. IPG achieves the highest fidelity circuit, approaching fidelity greater than $\mathcal{F}=1-10^{-7}$. L-BFGS is trapped in a local minimum. Note that while there were variations depending on the target unitary $\hat{C}$ and the initial guess, these trends were found to be quite representative of the observed behavior during optimization. The optimized angle for the $XX(\theta)\approx\mathbf{I}$ between $|S_{i,\sigma}\rangle$ and $|S_{i+1,\sigma}\rangle$, allowing that gate to be removed. c) Final pre-conditioner matrix $K$ from IPG, which estimates the inverse Hessian. Matrix elements concentrated around the main diagonal indicate a well conditioned estimate of the the inverse Hessian and a stable optimization result.
VI Error Mitigation
@@
In this section we outline two main error mitigation strategies employed on our circuit executions: Compiler lever (active) error mitigation strategies, and post-circuit execution error mitigation by exploiting knowledge of extensive symmetries of the original Fermi-Hubbard Hamiltonian and of the $\mathbb{Z}_{2}$ encoding a priori.
VI.1 Debiasing and Sharpening
@@
VI.2 Error mitigation based on global spin and local charge conservation
@@
The Fermi-Hubbard model intrinsically possesses global symmetries: the total number of fermions of each spin $\sum_{j}\hat{n}_{j\sigma}$ individually commutes with the Hamiltonian in Eq. (1), $\left[\sum_{j}\hat{n}_{j\uparrow},\hat{H}\right]=\left[\sum_{j}\hat{n}_{j\downarrow},\hat{H}\right]=0$. After the Jordan-Wigner transformation, this implies conservation of the total spin on each of the two sets of qubits on alternating sites (excluding the bond qubits), $\sum_{j}S^{z}_{j\sigma}$. Additionally, the lattice gauge theory encoding introduces ancillary bond qubits which must obey local charge conservation similar to quantum link models [22, 13, 23]: $\hat{Q}_{j}=(-)^{\hat{S}_{j\uparrow}+\hat{S}_{j\downarrow}}\,\hat{\tau}^{x}_{j-1,j}\,\hat{\tau}^{x}_{j,j+1}$. These conserved symmetries restrict the state of allowed spaces in the Hilbert space of $3N$ qubits for an $N$ site lattice based on the initial input state at $t=0$. In practice, due to gate infidelities and Trotter errors, the output of the quantum circuit does indeed have nonzero probabilities of going outside the conserved subspace. These states can be removed through post selection of the output histogram to restore the symmetries. In both the noisy statevector simulations and the QPU runs, we observed that the enforcing of global spin conservation in the up-spin and down-spin sector individually affected the observed quench dynamics significantly. As discussed in the next section and in Table 1, a little over 50% of shots were discarded in most QPU runs for the 4-site lattice, which results in a large fraction of the results surviving post-selection as opposed to what is often observed on current quantum hardware where more than 90% of the runs have to be discarded for deep circuits [35, 68].
VII QPU Results

### v2

@@ IV.1 Determining simplifiable subroutines
To optimize $\hat{C}$, we create a parametrizable circuit (in number of gates and overall depth) $\hat{C}_{A}$, known as an ansatz, that satisfies $\hat{C}_{A}\approx\hat{C}$. Our ansatz is shown in Fig. 2, and is comprised of single qubit $R_{X}(\theta)$ and $R_{Y}{\theta}$ gates, and a two qubit $XX(\theta)$ gate.
@@
In each iteration of gradient-based optimization (GBO), gate angles are generated by an optimizer. Then, using the matrix representation of $\hat{C}$ and $\hat{C}_{A}$, the similarity between the two circuits is given by $\mathcal{F}$, where
@@
Converging to a solution is limited by local minima, barren plateaus, initial conditions, noise when computing the gradient, and the number of tunable parameters. A larger depth $l$ results in more parameters to optimize over and the introduction of further minima and plateaus in the optimization problem. Therefore, an optimizer that is consistent in performance and resilient to these challenges is critical to finding gate angles that satisfy $\hat{C}\approx\hat{C}_{A}$.
Accelerated gradient descent techniques are needed to find suitable gate angles that satisfy $\hat{C}\approx\hat{C_{A}}$. This is done by minimizing a cost function $f(x_{*})=1-\mathcal{F}(x_{*})$, where $x_{*}$ is the gate angle vector. Formally, the goal is to compute a gate-angle vector $x_{*}\in\mathbb{R}^{d}$ such that
@@
Newton’s method [50] offers a method of accelerating convergence by using the Hessian of $f$. Specifically, Newton’s method uses preconditioning, which pre-multiplies the gradient with the inverse Hessian matrix at every iteration, resulting in local quadratic convergence rates [50] for strongly convex objectives. Despite its faster convergence rate, there are several issues in Newton’s method. (i) For empirical risk minimization, the per-iteration computational cost of Newton’s Method is $\text{O}(nd^{2}+d^{3})$. (ii) Secondly, the accelerated convergence of Newton’s method is guaranteed only if $f$ is strongly convex, an unlikely condition in quantum circuit optimization. (iii) It involves computing a matrix inverse at every iteration, which is highly unstable against process noise, such as qubit decoherence and gate infidelity. While Newton’s method offers superior convergence rates, these issues ultimately mean it is ill-suited for use in this setting.
IV.2 Iteratively Pre-conditioned Gradient Descent
We present iteratively pre-conditioned gradient descent (IPG) [35, 51], an accelerated gradient algorithm, to find suitable gate angles to surmount these challenges. IPG is a quasi-Newton method that iteratively estimates the inverse Hessian in a way that is more robust to process noise when compared to the explicit inverse Hessian and can remain stable even with a nonconvex cost. IPG achieves this by finding an approximate inverse Hessian with offset eigenvalues to help maintain stability, rather than explicitly taking the inverse Hessian like Newton’s method. One notable difference with other quasi-Newton methods, such as BFGS [50], which also construct an approximation of inverse Hessian, is that the preconditioner in IPG is not required to be symmetric and positive definite, a condition likely to be violated when implemented with noisy quantum gates.
We have previously shown the utility of the IPG algorithm on quantum circuit optimization [52] whereby faster convergence was observed over other optimizers like gradient descent and Adam [53]. Nevertheless, the exponential growth of the Hilbert space dimension with the number of qubits makes the problem of optimizing the entire quantum circuit simultaneously via IPG infeasible. Instead, we identified a 3-qubit unitary operator $\hat{C}$ which repeatedly appears at each Trotter step and optimize this subcircuit. By keeping the same timestep for each trotter step, the size of the circuit to be optimized stays fixed independent of the number of sites $N$ or qubit count $3N$. This leverages the model’s translational symmetry discussed in Sec. III.1 and the all-to-all connectivity of the trapped-ion QPU which makes implementing periodic boundary conditions connecting $i=1$ to $i=N$ straightforward. Further details of IPG are discussed in the Appendix, Sec. VIII.2.
IV.3 Optimizing the Ansatz
We optimize the gate angles in $\hat{C}_{A}$ at a depth of $l=3$ layers (Fig. 2a) using IPG, and benchmark IPG performance to L-BFGS and Adam, two optimizers commonly used for quantum circuit optimization. Cost is defined as 11-$\mathcal{F}$, where lower cost indicates better agreement between $\hat{C}_{A}$ and $\hat{C}$.
@@ IV.3 Optimizing the Ansatz
[…]
@@
After constructing the overall Trotter step using $\hat{C}_{A}$ with angles converged to by IPG, we verify that the reconstructed trotter step is the same as the original, using the convergence method outlined above. The average difference between states generated by a single parameterized vs directly executed trotter step is of order $10^{-4}$, sufficiently low to consider both circuits approximately equal.
@@
Figure 2: Gradient based optimization for circuit optimization. a) Modified ansatz $\hat{C}_{A}$ from [31] used for gradient based optimization to parametrize $\hat{C}$. Each layer is comprised of $R_{X}(\theta_{1})$, $R_{y}(\theta_{2})$, $XX(\theta_{3})$ on qubits $S_{i,\sigma}$, $B_{i,i+1}$, and $S_{i+1,\sigma}$ respectively. Three layers are used with a final $R_{X}(\theta)$ and $R_{Y}(\theta)$ rotation on each qubit. b) Result of GBO for 30 parameters using IPG, Adam, and L-BFGS. Best cost history over three trials and 128 iterations for each optimizer. IPG achieves the highest fidelity circuit, approaching fidelity greater than $\mathcal{F}=1-10^{-7}$. L-BFGS is trapped in a local minimum. Note that while there were variations depending on the target unitary $\hat{C}$ and the initial guess, these trends were found to be quite representative of the observed behavior during optimization. The optimized angle for the $XX(\theta)\approx\mathbf{I}$ between $|S_{i,\sigma}\rangle$ and $|S_{i+1,\sigma}\rangle$, allowing that gate to be removed. c) Final pre-conditioner matrix $K$ from IPG, which estimates the inverse Hessian. Matrix elements concentrated around the main diagonal indicate a well conditioned estimate of the the inverse Hessian and a stable optimization result.
IV.4 Decreasing two-qubit gatecount with von Neumann entropy
In our above optimization, the number of layers ($l$) in $\hat{C}_{A}$ is a preselected hyperparameter. Larger $l$ corresponds to a more expressive ansatz, at the cost of more gates. Ideally, we would like to minimize $l$ prior to applying IPG on $\hat{C}_{A}$.
To do so, we first modify the ansatz in Fig. 2 by adding an additional XX gate between $|S_{i,\sigma}\rangle$ and $|S_{i+1,\sigma}\rangle$. This creates the ansatz in Fig 3, ensuring both full single-qubit rotations for $l\geq 2$ and possible entanglement between each qubit in each layer. Now, the number and placement of entangling gates are the limiting factors to determine whether $\hat{C}_{A}$ can express $\hat{C}$. By evaluating the von Neumann entropy (VNE) of different permutations of two-qubit gate placements in our ansatz, we can predict the smallest two-qubit gate depth necessary to satisfy $\hat{C}_{A}\approx\hat{C}$.
IV.4.1 Procedure for predicting the smallest depth of an ansatz
$S(\rho_{i})=-Tr(\rho_{i}\log(\rho_{i}))$ (14)
Figure 3: Combining von Neumann Entropy (VNE) optimization with gradient based optimization for further two-qubit gate reduction on each subcircuit. a) Ansatz for the VNE optimization method. The ansatz number of layers is first truncated through the VNE method. Following this, gradient based optimization (GBO) is applied to find parameters such that $\hat{C}_{A}\approx\hat{C}$.During GBO, a number of single and two-qubit gate rotations go to $\mathcal{I}$, allowing further compression of the ansatz. b) Result of GBO over 13 parameters using IPG, Adam, and L-BFGS. Best cost history over three trials and 128 iterations for each optimizer. IPG achieves the highest fidelity circuit, approaching fidelity of $\mathcal{F}=1-10^{-7}$, with L-BFGS converging faster but to a lower fidelity (higher cost). c) Final pre-conditioner matrix $K$ from IPG, which estimates the inverse Hessian, similar to Fig. 2.
IV.4.2 Outcome of VNE optimization
With this method, $l\geq 3$ satisfied this expressibility criteria. Gradient-based optimization only converged to a solution at $l=4$, with the gates shown in Fig. 3 going to $\mathcal{I}$. The optimization sequence amounted to a thirteen 1Q gate reduction compared to $\hat{C}_{A}$, and more importantly, two fewer 2Q gates than $\hat{C}$. By also applying this process to $\hat{B}$, we were able to reduce $\hat{B}$ to an $XX(\theta)$ gate on $|\tau_{i-1,i}\rangle,|\tau_{i,i+1}\rangle$ and a single $R_{y}(\theta)$ on $|\tau_{i-1,j}\rangle$.
Table 1 summarizes the overall reduction in two-qubit gates per Trotter step, with an overall reduction from 14N to 9N. This decrease is substantial: Assuming 400 2Q gate executions as the Aria-1 algorithmic qubit limit [54], we are able to increase the number of faithful Trotter steps simulated on the Aria-1 QPU by $\approx 50\%$, from 4 to 7 for $N=6$.
Table 1: 2-qubit gate (2Q) cost per trotter step for modelling NN sites of the FHM encoded as a $\mathbb{Z}_{2}\$LGT. Results summarized for a) the initial direct trotter step decomposition b) by parametrizing $\hat{C}$ and applying gradient-based optimization (GBO) on $\hat{C}_{A}$, and c) through applying GBO with the non Neumann Entropy (VNE) method on $\hat{C}_{A}$ and a parametrized $\hat{B_{A}}$. Our final method is a 36% reduction in 2Q cost per trotter step without comprising on exactness of the $\mathbb{Z}_{2}\$LGT FHM encoding.
Trotter Step Variant 2Q Cost
Direct 14N
GBO on $\hat{C}_{A}$ 12N
GBO and VNE on $\hat{C}_{A}$ and $\hat{B}_{A}$ 9N
V Error Mitigation
@@
In this section we outline two main error mitigation strategies employed on our circuit executions: Compiler lever (active) error mitigation strategies, and post-circuit execution error mitigation by exploiting knowledge of extensive symmetries of the original Fermi-Hubbard Hamiltonian and of the $\mathbb{Z}_{2}\$LGT encoding a priori.
V.1 Debiasing and Sharpening
@@
V.2 Error mitigation based on global spin and local charge conservation
@@
The Fermi-Hubbard model intrinsically possesses global symmetries: the total number of fermions of each spin $\sum_{j}\hat{n}_{j\sigma}$ individually commutes with the Hamiltonian in Eq. (1), $\left[\sum_{j}\hat{n}_{j\uparrow},\hat{H}\right]=\left[\sum_{j}\hat{n}_{j\downarrow},\hat{H}\right]=0$. After the Jordan-Wigner transformation, this implies conservation of the total spin on each of the two sets of qubits on alternating sites (excluding the bond qubits), $\sum_{j}S^{z}_{j\sigma}$. Additionally, the lattice gauge theory encoding introduces ancillary bond qubits which must obey local charge conservation similar to quantum link models [23, 13, 24]: $\hat{Q}_{j}=(-)^{\hat{S}_{j\uparrow}+\hat{S}_{j\downarrow}}\,\hat{\tau}^{x}_{j-1,j}\,\hat{\tau}^{x}_{j,j+1}$. These conserved symmetries restrict the state of allowed spaces in the Hilbert space of $3N$ qubits for an $N$ site lattice based on the initial input state at $t=0$. In practice, due to gate infidelities and Trotter errors, the output of the quantum circuit does indeed have nonzero probabilities of going outside the conserved subspace. These states can be removed through post selection of the output histogram to restore the symmetries. In both the noisy statevector simulations and the QPU runs, we observed that the enforcing of global spin conservation in the up-spin and down-spin sector individually affected the observed quench dynamics significantly. As discussed in the next section and in Table 1, about 50% of shots were discarded in most QPU runs for the 6-site lattice, which results in a large fraction of the results surviving post-selection as opposed to what is often observed on current quantum hardware where more than 90% of the runs have to be discarded for deep circuits [36, 58].
VI QPU and Noisy Statevector Simulator Results